A theater sold $330 worth of tickets. An adult ticket costs $2.00 each and a child’s ticket cost $1 each. If total tickets sold were 190, how many tickets of each kind were sold?
Answer by checkley77(12844)

A+C=190
A=190-C
2A+1C=330
2(190-C)+C=330
380-2C+C=330
-C=330-380
-C=-50
C=50 ANS. FOR THE NUMBER OF CHILDREN TICKETS.
A=190-50=140 ANS. FOR THE NUMBER OF ADULT TICKETS.
PROOF:
2*140+50=330
280+50=330
330=330